### v2

# VitaGlyph: Vitalizing Artistic Typography with Flexible Dual-branch Diffusion Models

Abstract
Artistic typography is a technique to visualize the meaning of input character in an imaginable and readable manner. With powerful text-to-image diffusion models, existing methods directly design the overall geometry and texture of input character, making it challenging to ensure both creativity and legibility. In this paper, we introduce a dual-branch, training-free method called VitaGlyph, enabling flexible artistic typography with controllable geometry changes while maintaining the readability. The key insight of VitaGlyph is to treat input character as a scene composed of a Subject and its Surrounding, which are rendered with varying degrees of geometric transformation. To enhance the visual appeal and creativity of the generated artistic typography, the subject flexibly expresses the essential concept of the input character, while the surrounding enriches relevant background without altering the shape, thus maintaining overall readability. Specifically, we implement VitaGlyph through a three-phase framework: (i) Knowledge Acquisition leverages large language models to design text descriptions for the subject and surrounding. (ii) Regional Decomposition detects the part that most closely matches the subject description and divides the input glyph image into subject and surrounding regions. (iii) Typography Stylization first refines the structure of the subject region via Semantic Typography, and then separately renders the textures of the Subject and Surrounding regions through Controllable Compositional Generation. Experimental results demonstrate that VitaGlyph not only achieves better artistry and readability but also manages to depict multiple customized concepts, facilitating more creative and pleasing artistic typography generation. Our code will be made publicly available.
Figure 1: Examples of our VitaGlyph. The rendered images not only convey the intrinsic semantics of input characters (i.e., first row) but also visualize the user-specific concepts (i.e., second row).
1 Introduction
Figure 2: Overview of VitaGlyph. VitaGlyph divides the input glyph image into Subject and Surrounding and follows a three-phase pipeline: (i) Given an input character, Knowledge Acquisition leverages large language models to design prompts for both subject and surrounding. (ii) With the subject prompt and glyph image, Regional decomposition detects the part that best matches the subject description and splits the glyph into subject and surrounding regions. (iii) Typography Stylization first transforms the Subject structure via Semantic Typography and then separately renders textures for both regions through Controllable Compositional Generation.
Artistic typography focuses on emphasizing the visual meaning of words while maintaining their legibility. This technique has widespread applications across various fields such as commercial advertising [10, 22], social education [43, 9], and design promotion [52, 45]. Traditional artistic typography involves manually applying design principles such as color, spacing, and font choice to convey meaning [11, 17]. However, this process is often time-consuming and lacks the flexibility to dynamically adapt to varying content.
Recent works [17, 11, 31] leverage the powerful capabilities of pre-trained text-to-image diffusion models [40] to advance the design process. Despite significantly improving productivity and diversity, it remains challenging to flexibly portraying the visual characteristics (e.g., the transformation of texture and geometry) of input words without compromising readability. FontStudio [31] focuses solely on depicting the given word through style and texture while neglecting geometric deformation. In contrast, Word-as-image [17] and Khattat [16] target black-and-white designs by altering the shape of selected letters. WordArt Designer [11] attempts to integrate various transformations into semantic typography but may sacrifice readability when illustrating concepts that involve visible changes to the overall glyph geometry. For example, in Fig.5, the “tomato” resembles several tomatoes more than the character itself. MetaDesigner[12], as an integrated system employing multiple models, requires substantial manual effort and extensive engineering techniques to achieve promising artistic typography. Moreover, existing methods can express only a single interpretation of a given concept, further limiting flexibility and user control over the technique.
In this work, we introduce VitaGlyph to facilitate flexible and imaginative artistic typography with controllable transformations of glyph geometry while maintaining readability. The core idea of VitaGlyph is to divide an input character (e.g., “rose” in Fig. 2) into the Subject (e.g., “red rose”) and the Surrounding (e.g., “brown vignette, leaf, slender branch…”). The Subject expresses the essential concept of the input character, while the Surrounding enriches the relevant background without altering the shape.
We implement VitaGlyph through a three-phase approach: Knowledge Acquisition, Regional Decomposition, and Typography Stylization. In the first phase, we leverage large language models (e.g., ChatGPT [32]) to acquire prior knowledge of the input character, designing text prompts for the Subject and Surrounding. In the second phase, Grounding-DINO [23] is used to identify the part of the glyph that structurally resembles the subject concept, allowing the glyph image to be decomposed into subject and surrounding regions.
During the third phase, we render the subject and surrounding regions with their corresponding prompts through Controllable Compositional Generation, which we propose inventively. To better align with the intrinsic semantics of the subject concept, Semantic Typography is used to deform the geometry of the subject image before the rendering process. Notably, VitaGlyph enables users to incorporate multiple customized concepts into a single glyph, thereby enhancing user control and fostering greater creativity.
Extensive experimental results demonstrate that the proposed VitaGlyph offers significant advantages in character transformations while maintaining strong readability. Furthermore, by employing independent diffusion models, VitaGlyph enables artistic text generation with multiple customized concepts—an accomplishment that other approaches fail to achieve. Additionally, the introduction of controllable conditions provides users with greater flexibility in designing personalized content such as signatures and advertisements, as shown in Fig. 6.
To conclude, we summarize our main contributions as follows:
  • We introduce the training-free VitaGlyph, which enables flexible and imaginable artistic typography through controllable transformations of glyph geometry, while maintaining readability by dividing input characters into Subject and Surrounding.
  • VitaGlyph portrays the Subject as the essential concept of the input character and uses the surrounding to enrich the background without harming readability.
  • Comprehensive experiments on collected glyph images demonstrate that VitaGlyph surpasses competitors in both creativity and readability.
  • VitaGlyph can be flexibly extended to depict multiple subject concepts, allowing users to add customized descriptions while maintaining intrinsic semantics.
2 Related Work
2.1 Artistic Text Generation
Artistic text generation [3] is divided into text stylization, which focuses on visual effects [44, 27], and Semantic Typography, which deforms text to match its meaning [42, 11, 16]. Models like eDiff-I[4], DeepFloyd [19], and TextDiffuser-2 [7] enhance generation using LLMs, while others [46, 26, 37] add motion.
However, these methods often capture only one layer of meaning. We propose a dual-branch model that combines stylization and typography for more complex text generation.
2.2 Large Language Models for Image Generation
Large language Models(LLMs) [1] have extend their impressive abilities from language tasks to vision synthesis. Most generation-related works [21, 47, 11] use LLMs for prompt understanding, layout planning and typography design. Due to these insights, we incorporate the LLM to design artistic characters.
2.3 Controllable and Compositional Generation
ControlNet-based methods [30, 49, 15, 24, 51] improve image synthesis control with conditions like segmentation or depth maps, and some studies [34, 48, 31] apply these to text generation. Leveraging the geometric nature of glyphs, we use ControlNet with glyph conditions as our foundation.
For multi-concept synthesis, works [6, 2, 29, 39] optimize noise maps for concept co-occurrence, while [18] uses Diffusion Models with LoRAs [14], and [25] integrates residual embeddings and layout priors.
Few methods [50] combine compositional generation with controllable conditions. Our work achieves pixel-level control in compositional generation using ControlNet and user-specific masks.
3 Method
Artistic typography generation task aims to visually convey the meaning of input characters in an appealing and comprehensible way while also ensuring legibility. To this end, we introduce a dual-branch method, namely VitaGlyph, which performs adaptive rendering on the Subject and Surrounding components of input glyph. The Subject undergoes structural modifications to flexibly express the intrinsic semantics of the input character, reflecting strong artistic essence. Meanwhile, the Surrounding enriches the details of the subject while maintaining overall readability by adhering to the character’s structure throughout the generation process.
As illustrated in Fig. 2, VitaGlyph consists of three phases: Knowledge Acquisition, Regional Decomposition, and Typography Stylization. In the Knowledge Acquisition phase, we ask the large language models (e.g., ChatGPT [32]) to acquire the prior knowledge of the Subject and Surrounding and design their text prompts. In Regional Decomposition phase, we parameterize the input character into glyph image, followed by employing Grounding-DINO (i.e., GDINO) to detect the subject part. Finally, during Typography Stylization phase, we firstly transform the subject’s structure and then render subject and surrounding images with their corresponding prompts through Controllable Compositional Generation (CCG).
3.1 Knowledge Acquisition
The goal of Knowledge Acquisition is to convert input characters into descriptive language prompts for subjects and surroundings that are interpretable by text-to-image diffusion models. Manually designing these prompts or relying solely on the input characters themselves (e.g., “rose”) can be challenging, as these methods are either time-consuming or produce descriptions that are difficult to interpret, particularly for abstract words like “win” or “lose” [11].
The large language models like ChatGPT [32] have demonstrated significant capabilities in understanding, imagining, and interpreting. Following prior work [11], we utilize ChatGPT [32] to generate concrete and detailed text descriptions for input characters. However, while [11] primarily focuses on depicting the basic concept of an input character, our approach aims to visualize it more completely and vividly as a compositioanl scene (i.e., subject and its surroundings). Specifically, given a query prompt $Q_{c}$ such as “Help me design an artistic character for …”, we employ ChatGPT to interpret the input character $c$ into a subject prompt $P_{\mathrm{sub}}$ and a surrounding prompt $P_{\mathrm{surr}}$ as follows:
$P_{\mathrm{sub}},P_{\mathrm{surr}}=\texttt{GPT}(Q_{c},c),$ (1)
Supplementary materials provide more detailed examples.
3.2 Regional Decomposition
Given subject description $P_{\mathrm{sub}}$, the Regional Decomposition phase aims to convert the input character $c$ into glyph image $I$ and then divide the image $I$ into subject and surrounding regions at the image level. The subject region corresponds to the part of the image that is most structurally similar to the essential concept described in $P_{\mathrm{sub}}$.
3.2.1 Parameterization
Given an input character $c$, we use the FreeType font library [8] to extract the outline of the character and convert it into cubic Bézier curves. After that, the Differentiable Vector Graphics(DiffVG) [20] is utilized to raster obtained Bézier curves into get a final glyph image $I$.
3.2.2 Region Interpretation
Figure 3: Results of Grounding DINO with filtering and ranking strategies. The red box shows the detected subject region $I_{\mathrm{sub}}$ that resembles $P_{\mathrm{sub}}$ most structurally.
To preserve the shape of input glyph (i.e., readability) as much as possible, we choose the region most similar to the structure of subject concept as subject region (e.g., “snow” in Fig. 3). Compared to previous works that use random region selection [11, 12] or manual region selection [17], we leverage Grounding-DINO [23] to detect the part of input image $I$ that matches the subject description $P_{\mathrm{sub}}$, thereby efficiently bridging the semantic gap between the selected regions and the subjects, while significantly reducing time. As the detected results contain multiple bounding boxes of varying sizes, we filter out those with unreasonable sizes and any that fall below the confidence threshold [^1] We set the reasonable proportion of the bounding box area relative to the image area between 0.4 and 0.6, and the confidence threshold to 0.5 Then we select the bounding box with the highest confidence score as the subject region:
$I_{\mathrm{sub}},M=\texttt{GDINO}_{\mathrm{rank}}(I,P_{\mathrm{sub}}),$ (2)
where $I_{\mathrm{sub}}$, $\texttt{GDINO}_{\mathrm{rank}}$ and $M$ denote the subject image, Grounding-DINO with filtering and ranking strategies and the mask area of the Subject, respectively. As shown in Fig. 3, due to Grounding-DINO’s open-domain detection capabilities, the $I_{\mathrm{sub}}$(red bounding box) of characters that best resemble the $P_{\mathrm{sub}}$ can be easily detected. There are no specific structural requirements for the surrounding image $I_{\mathrm{surr}}$, which is expected to maintain legibility. Thus, we consider the remaining part of the glyph image as the surrounding region:
$I_{\mathrm{surr}}=I-I_{\mathrm{sub}}.$ (3)
3.3 Typography Stylization
As mentioned above, the subject and surrounding components play different roles in depicting the visual semantics of input characters, so their structures and textures are rendered in different ways. Firstly, the structure of subject image is flexibly altered and enriched by Semantic Typography, while the surrounding remains unchanged. Secondly, we render their texture style through two independent ControlNets (i.e., as $\mathrm{ControlNet}_{\mathrm{sub}}$ and $\mathrm{ControlNet}_{\mathrm{surr}}$) in a compositional manner, which we term as Controllable Compositional Generation (CCG).
3.3.1 Semantic Typography
Previous works [17, 11] rely on time-consuming score distillation sampling loss [36] to transform the structure of input glyphs. To improve efficiency, we directly utilize the depth-to-image model in LDM [40] to adjust the geometry of $I_{\mathrm{sub}}$. Specifically, we adopt the SDEdit algorithm [28] to perform geometry transformation, conditioned on the depth map $I_{\mathrm{sub}}$ and subject prompt $P_{\mathrm{sub}}$. It can be described as follows:
$I^{*}_{\mathrm{sub}}=\texttt{SemTypo}(P_{\mathrm{sub}},I_{\mathrm{sub}}).$ (4)
3.3.2 Controllable Compositional Generation
Figure 4: Overview of Controllable Compositional Generation. The Controllable Compositional Generation utilizes Stable Diffusion and two independent ControlNets to synthesize and fuse the subject and surrounding regions with the given mask $M$.
Figure 5: Visualization of artistic typography results compared with existing methods. WordArt Designer and MetaDesigner apply transformations to entire characters, often sacrificing readability, as seen in “mushroom”, where the bottom becomes unrecognizable. Additionally, some characters blend into the background without independent diffusion modeling, such as “cabinet” in WordArt and “tomato” in MetaDesigner. In contrast, our method preserves better readability with controllable transformations, separates characters from the background and maintains a visually engaging design. We also provide a comparison with Word-as-image in the English version.
To render subject and surrounding images with their corresponding prompts independently, we apply two ControlNets to separately transfer texture style in a compositional manner. For the subject image $I_{\mathrm{sub}}$, after transforming its structure in SemTypo into $I^{*}_{\mathrm{sub}}$, we use the segmentation version of ControlNet, referred to as $\mathrm{ControlNet}_{\mathrm{sub}}$. For the surrounding image $I_{\mathrm{surr}}$, we select the scribble version of ControlNet, denoted as $\mathrm{ControlNet}_{\mathrm{surr}}$ to maintain the character layout while adding creativity and artistic flair. This dual-branch diffusion process decouples the rendering of subject and surrounding, allowing the generated artistic typography to incorporate different concepts.
Specifically, we first encode $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$ into text embeddings $\boldsymbol{p}_{\mathrm{sub}}$ and $\boldsymbol{p}_{\mathrm{surr}}$, respectively. Both $\mathrm{ControlNet}_{\mathrm{sub}}$ and $\mathrm{ControlNet}_{\mathrm{surr}}$ share the same random noise $\epsilon_{T}\in\mathcal{N}(0,I)$ as the initial latent, conditioned on their corresponding structures and text embeddings. We use the DDIM algorithm [41] for sampling. To fuse subject and surrounding textures at timestep $t$, given subject mask $M$, we combine the predicted noises $\epsilon^{\mathrm{sub}}_{t}$ and $\epsilon^{\mathrm{surr}}_{t}$ to compute the overall noise $\epsilon^{\mathrm{overall}}_{t}$ as:
$\epsilon^{\mathrm{overall}}_{t}=\gamma\cdot M\cdot\epsilon^{\mathrm{sub}}_{t}+(1-M)\cdot\epsilon^{\mathrm{surr}}_{t},$ (5)
where $\gamma\in(0,1)$ is the hyper-parameter to adjust the prominence of the Subject within the overall image.
The overall Controllable Compositional Generation process is illustrated in Fig. 4. Since the sampling processes for the subject and surrounding are conducted independently, and the U-Net outputs are independent as well, directly merging them can result in a visually incoherent image, as seen in Fig.8 (b). In our method, we resolve this fragmentation issue by allowing the subject and surrounding diffusion processes to share the same unconditional noise $\epsilon^{\mathrm{uc}}_{t}$ during Controllable Compositional Generation. $\epsilon^{\mathrm{overall}}_{t}$ can be further harmonized as follows:
$\hat{\epsilon}^{\mathrm{overall}}_{t}=\epsilon^{\mathrm{uc}}_{t}+s(\epsilon^{\mathrm{overall}}_{\mathrm{t}}-\epsilon^{\mathrm{uc}}_{t}),$ (6)
where $s$ represents the guidance scale.
4 Experiments
Figure 6: Visualization of VitaGlyph’s applications. The experiments show that our method can be applied to other languages and fonts. Additionally, VitaGlyph can be extended to generate with multiple customized concepts.
Methods Language CLIP Score $\uparrow$ GPT-4o(TIM) $\uparrow$ GPT-4o(OCR) $\uparrow$ FID Score $\downarrow$
WordArt Designer [11] Chinese 23.93 94.98 88.67 249.12
VitaGlyph 25.40 96.38 92.89 239.27
Word-as-image [17] English 29.01 71.17 80.52 255.28
VitaGlyph 29.16 75.98 81.96 247.17
Table 1: Quantitative results. Metrics include CLIP score, GPT-4o Text-Image Matching(TIM), GPT-4o Optical Character Recognition(OCR) and FID Score. Our model outperforms WordArt Designer and Word-as-image by a significant margin.
4.1 Implementation Details
4.1.1 Benchmark
We collect two hundred Chinese words (including single and multiple characters) as an evaluation benchmark and ask ChatGPT to annotate subject and surrounding descriptions. Considering that Word-as-image [17] only generates English words, we also translate the collected benchmark into English. More benchmark details are provided in the supplementary materials.
4.1.2 Evaluation Metrics
We use CLIP [38] image-text similarity and FID [13] scores to assess prompt alignment and readability. To better align with human evaluators, we also employ GPT-4o [33] to determine whether the rendered images match the text descriptions (GPT-4o Text-Image Matching) and can be correctly recognized as input characters (GPT-4o OCR).
4.2 Quantitative Comparisons
As shown in Tab. 1, we conduct quantitative experiments on the benchmark mentioned above. For comparisons, we leverage WordArt Designer [11] for Chinese words and Word-as-image [17] for English version. Since Word-as-image can only generate black text with white backgrounds, we apply ControlNets on it for texture rendering, as done in prior work [11]. Since MetaDesigner has not open-sourced the code and we can only generate results one by one through the website demo, we have only conducted a qualitative comparison with it.
Our VitaGlyph outperforms WordArt Designer in all metrics, especially in OCR with a 4.22% improvement. We believe this is because, with the controllable and proper transformation of glyph geometry, our method maintains promising readability. Meanwhile, for English words, VitaGlyph also demonstrates strong generalization capabilities, especially in TIM with a 4.81% improvement. This further proves that our method can better express essential concept with elaborately designed subject transformations.
Since all methods’ output images are in color while the ground-truth images for FID evaluation are black and white, all methods result in high FID scores. However, we still outperform other baselines.
4.3 Qualitative Comparisons
We compare VitaGlyph with WordArt Designer and MetaDesigner, both of which are strong baselines for Chinese artistic typography. We also design English words for a qualitative comparison with Word-as-image. The input prompts of WordArt Designer, MetaDesigner and Word-as-image consist of the concatenation of the subject descriptions and the surrounding descriptions. The input prompts for the English word is manually designed to include Nashville’s specialties, such as “country music, guitar, fruit” and so on.
In Fig. 5, we see that VitaGlyph renders input characters with higher quality and legibility than WordArt Designer and achieves results comparable to, or even better than, MetaDesigner. For example, in column $1$, VitaGlyph visualizes the concept of “tomato” more concisely, incorporating related elements like “fries.” However, the image from WordArt Designer suffers from structural damage, and that of MetaDesigner blends the subject and background, visibly harming readability. As for Word-as-Image, it excels advantage in structural transformation but lacks artistic flair as it can only generate black-and-white typography. In contrast, VitaGlyph properly generates flexible and imaginable artistic typography with the controllable transformation. More examples generated by our VitaGlyph can be found in Fig. 1 and supplementary materials.
4.4 User Study
Methods Text Align. Char. Align. Artistic
WordArt Designer 29.17% 31.67% 37.50%
Word-as-image 25.83% 32.50% 19.17%
VitaGlyph 45.00% 35.83% 43.33%
Table 2: User Preference Study. Thirty users participated in the survey, selecting images that best represent the concept (Text Align.), match the character (Char. Align.), and show the highest artistic mastery (Artistic).
We conduct a user study to evaluate our method against WordArt Designer and Word-as-Image. For customized multi-concept generation, users are shown the generated artistic text from each method along with ten reference images of the character. They are then asked to select the image that most represents the concept, the image that best matches the character itself, and the image that is most aesthetically pleasing. The questionnaire contains 80 sets for each question, and 30 responses were collected in total. The statistical results in Tab. 2 show that users prefer the rendered images from VitaGlyph across all metrics.
4.5 Ablation Studies
[…]
TIM $\uparrow$ OCR $\uparrow$
w/o $\texttt{GDINO}_{\mathrm{rank}}$ 23.90 94.95 89.15 255.34
w/o SemTypo 23.53 95.27 92.98 241.58
w/o shared-$\epsilon_{T}$ 23.82 95.22 90.38 244.77
Ours 25.40 96.38 92.89 239.27
Table 3: Ablation of components in our method. We provide our ablation studies’ quantitative results with different modules removed.
4.5.1 Effect of Region Interpretation
Figure 7: Visualization of ablation study on Region Interpretation. (a) Regard the entire image as the subject region. (b) Randomly select a part of the image as the subject region (i.e., in red box). (c) Employ $\texttt{GDINO}_{\mathrm{rank}}$ to select the most appropriate part as the subject region. Results of $I^{*}_{\mathrm{sub}}$s and generated images indicate that VitaGlyph with $\texttt{GDINO}_{\mathrm{rank}}$ is superior.
To study the effect of our Region Interpretation from Sec.3.2.2, we conduct ablation experiments without $\texttt{GDINO}_{\mathrm{rank}}$. Specifically, we regard $I$ as $I_{\mathrm{sub}}$, while keeping the settings for SemTypo and CCG unchanged. We solely utilize a $\mathrm{controlnet}_{\mathrm{sub}}$ with $I^{*}_{\mathrm{sub}}$ as the additional condition. The input prompt is a concatenation of $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$. The results are shown in Fig. 7 (a). Additionally, we also randomly select the subject region rather than utilize $\texttt{GDINO}_{\mathrm{rank}}$, with all other settings unchanged. The results are presented in Fig. 7 (b).
Fig. 7 shows that without Region Interpretation, VitaGlyph faces challenges: In Fig. 7 (a), the overall deformation of characters is abnormal and some key concepts, such as “fries” in “tomato” (3rd row), are missing. In Fig. 7 (b), unreasonable region selection makes it difficult to ensure both prompt alignment and readability simultaneously. For example, the structure of “lavender” is deformed (1st row) and the features of “cat” are not accurately generated (2nd row).
For quantitative results, comparing to rows 2 and 5 of Tab. 3, the CLIP score improves by 1.5%, and accuracy increases by 1.43% in GPT-4o(TIM) and 3.09% in GPT-4o(OCR). This is likely because Grounding-DINO identifies the part most resembling the subject prompt, making glyph geometry transformation easier and more natural.
4.5.2 Effect of Semantic Typography
Figure 8: Visualization of ablation studies on SemTypo and shared-$\boldsymbol{\epsilon_{T}}$. Input characters are shown at the top of the sub-figures. The different columns in sub-figures (a) and (b) display experimental results with/without SemTypo or shared-$\epsilon_{T}$.
Although Grounding-DINO identifies the part similar to the subject prompt $P_{\mathrm{sub}}$, the Semantic Typography is essential for structural transformation. Ablation experiments without this module show that images fail to capture the character’s essence or cause the Subject to appear fragmented in Fig. 8 (a). This is due to the gap between $I_{\mathrm{sub}}$ and real-world objects, which the Depth-to-image module helps bridge.
Based on the rows 3 and 5 from Tab. 3, we can conclude that VitaGlyph with Semantic Typography would enrich the semantic expression but sacrifice little readability.
4.5.3 Effect of shared-$\boldsymbol{\epsilon_{T}}$
As VitaGlyph employs two ControlNets to individually render subject and surrounding regions, we investigate the impact of shared initial noise (i.e., shared-$\epsilon_{T}$) on the final results. Fig. 8 (b) indicates that not using shared-$\epsilon_{T}$ leads to inconsistent appearance and unrealistic lighting. For example, in the 1st column of “rose” and “monkey”, the colors of subject and surrounding regions are visibly incoherent. Additionally, rows 4 and 5 of Tab. 3 show that using shared-$\epsilon_{T}$ improves prompt alignment and readability, consistent with visualization results.
4.5.4 The Strength of Added Noise in SemTypo
Figure 9: Ablation Study on D2I’s strength $\mathrm{DS}$ and ControlNet’s scale $\mathrm{CS}$. (a) shows that as $\mathrm{DS}$ increases, the deformation of $I^{*}_{\mathrm{sub}}$ grows during SemTypo. (b) demonstrates the generated image morphology aligns more closely with $I^{*}_{\mathrm{sub}}$ as $\mathrm{CS}$ increases.
Semantic Typography uses the SDEdit algorithm [28] based on depth-to-image diffusion model to adjust the geometry of subject region. As shown in Fig. 9 (a), as the strength of added noise ($\mathrm{DS}$) increases, more structural details are added to the subject region while remaining the black-and-white color scheme unchanged. When $\mathrm{DS}$ reaches $0.85$, the subject region contains enough details of the given concept, e.g., the ears and tails of cats.
4.5.5 The Conditioning Scale in ControlNet
We fix the $I^{*}_{\mathrm{sub}}$ and only adjust the $\mathrm{ControlNet}_{\mathrm{sub}}$’s scale $\mathrm{CS}$. From Fig. 9 (b), we observe that as the $\mathrm{CS}$ increases, the generated part of Subject becomes more clear and more similar to the $I^{*}_{\mathrm{sub}}$, shown in row 1. However, being too similar to the subject can lead to detail loss, as seen in the row 2 where the flower’s stamen disappears from $\mathrm{CS}=1.0$.
4.6 Applications on Fonts and Languages.
In first row of Fig. 6, we demonstrate VitaGlyph’s effectiveness across various fonts and languages, such as “Bold” in English, “Song Traditional” in Chinese and “Noto Sans” in Korean. VitaGlyph not only generates artistic glyph images but also preserves the original font styles.
Moreover, VitaGlyph extends from single to multiple customized subject concepts. As illustrated in Fig. 6, it designs subjects like “the Earth”, “compass” and “colorful doughnut” within the artistic typography of “Google”. This plays a crucial role in areas such as commercial advertising, brand identity, and corporate image design. More examples can be seen in Fig. 1 and supplementary materials.
5 Conclusion
This paper presents VitaGlyph, a novel approach to artistic typography that balances creativity and readability through controllable glyph transformation. By sdistinguishing between the subject and surrounding elements, VitaGlyph enhances visual representation while preserving the character’s core meaning. Our method makes significant contributions to artistic text generation. Future work will focus on further advancements in generative AI to enhance the adaptability of typography across diverse contexts and styles. VitaGlyph will significantly enrich artistic expression and enabling more personalized, dynamic designs.
F Benchmark Details
We collect two hundred Chinese words (including single and multiple characters) as evaluation benchmark, as shown in Fig .K. Every word we generate one hundred samples and we totally create twenty thousand images as a dataset. Considering that Word-as-image [17] only generates English words, we also translate the collected benchmark into an English version with fifty samples per word and ten thousand images as a dataset at all.
For words’ descriptions, we ask ChatGPT [32] to annotate subject and surrounding prompts. Since other method, Word-as-image [17], WordArt Designer [11] and MetaDesigner [12] can only leverage one input prompt, we concatenate the subject and surrounding prompts.
G Deployment Details
In phase Knowledge Acquisition, we use GPT3.5 as our Large Language Model and the output of the LLM is JSON format. In phase Regional Decomposition, we use Grounding-DINO with its hyper-parameters all default to the  [23]. After that, we select the bounding boxes with its confidence over 0.5 and its area proportion of the whole image between 0.4 and 0.6 and then choose the highest confidence as the subject part. In phase Typography Stylization, we use Stable Diffusion-1.5 as our base model. All generated images are $512\times 512$ for both Depth-to-image and CCG module. Though we haven’t used larger modules (e.g., SD-XL [35] and DALL-E 3 [5]) to generate higher quality images, our results still demonstrate a strong generative capability. It is easy to deployment our work with one Nvidia RTX 3060 12GB GPU, taking about 40 seconds to generate 1 image.
H Additional Results
We provide more Chinese artistic typography examples in Fig. J and examples with multiple customized concepts in Fig, L. In Fig. L, we can observe that we incorporate “coffee beans” and “doughnut” in artistic typography of “Starbucks”, with its texture style is “green and white background”. And we can also generate “Hello World” in Chinese, Japanese, Korean and English with concepts like “Robot” and “The Earth”.
I Limitations
Fig. L’s “Jordan” shows that our method would exhibit some deficiencies when generating humans. We believe this is due to the base model, StableDiffusion-1.5. In future work, we will address and refine these issues. Additionally, while we utilize multiple parallel diffusion models for compositional generation, incorporating more than two concepts with different models poses challenges in terms of memory usage and computation time and we will optimize our implementation as soon as possible.
J Future Work
In addition to addressing the limitations mentioned above, there exist several promising avenues for future research. [26, 46] delve into the animation of artistic words. Offering insights into the harmonious integration of movement and typography artistry, these studies pave the way for us to explore into the dynamic and kinetic dimensions of typographic expression. Additionally, Our approach are the first to engage artistic typography with compositional generation. We are honored to develop a artistic typography dataset and improve our benchmark with more methods and metrics. This provides a foundation for the development of innovative methods to enhance artistic typography techniques.
K Social Impacts
VitaGlyph has the potential to significantly impact the creative industry by offering new avenues for designers and artists to explore innovative and appealing typographic designs. This technology not only enhances artistic expression and readability but also has the capacity to elevate user experiences in fields such as advertising, branding, and digital media. Moreover, VitaGlyph’s interdisciplinary approach, combining text generation, image processing, and typography design, can foster cross-domain collaborations, driving technological advancements and fostering innovation. Without proper guidance and ethical considerations, individuals may engage in harmful practices such as image forgery or digital impersonation. We advocate for the development of legal frameworks that address AI-generated content, including penalties for malicious use.
Figure L: More examples of VitaGlyph’s application. Our method can be applied to various languages and fonts. With multiple user-specific concepts injected and the utilization of Controllable Compositional Generation, we manage to design logo or slogan like “CVPR”, “VitaGlyph” and “Hello World”..